# LIME: Localized Image Editing via Attention Regularization in Diffusion Models

### v1

@@ 8.2 Visual Comparison
Figure 8: VQGAN-CLIP [9] Comparison. CLIP-T metrics are reported below each image and calculated between the output caption and the corresponding image. Input images and edit instructions are pictured in the first column. Ground truth edit images are taken from the MagicBrush dataset.
Blended Latent Diffusion.
As shown in Tab. 2, Blended Latent Diffusion [3] has better results than baselines and our method. However, as shown in Fig. 9, even if their method can perform the desired edit on the given mask (RoI) from the user, (a) it distorts the location of the features, e.g., heads of the birds, and (b) it loses the information of the object in the input image and creates a new object in the RoI, e.g., blanket in (b). On the other hand, our method performs visually appealing edits on the input images considering the given edit instructions by preserving as many details from the input image as possible. This is also highlighted by a significantly lower Distance metric for our method in Tab. 2.
@@
Diffusion Disentanglement.
Wu et al. [48] propose a disentangled attribute editing method. Figure 10 shows edit types such as (a) texture editing and (b) replacing the object with a similar one. Diffusion Disentanglement on (a) alters the background objects in the image, e.g., adding snow on and changing the shape of the branch, and also changes the features of the object of interest, e.g., removing the tail of the bird. On (b), it fails to perform the desired edit altogether. Moreover, it requires a GPU with > 48 GB RAM [^2] https://github.com/UCSB-NLP-Chang/DiffusionDisentanglement, and one image takes approximately 10 minutes on an NVIDIA A100 80GB to generate the edited version. In comparison, our method achieves higher visual quality and takes 25 seconds to complete on NVIDIA A100 40GB with a GPU RAM usage of 25 GB.
Input Image
DiffusionDisent. [48]
Ours
Change the robin to a silver robin sculpture.
Turn the brown horse into a pink unicorn.
8.3 Qualitative comparison on segmentation maps
Our method proposes an alternative segmentation method based on the clustering of intermediate features of the diffusion process. In this section, we provide a qualitative comparison to other segmentation methods. LPM [33] uses self-attention features from one resolution, such as $32\times 32$, while our method leverages the intermediate features from different resolutions to enhance the segmentation map. Then, both apply a clustering method to find the segments in the input image. Another way to find segments is by using large segmentation models, e.g., SAM [22], ODISE [49] …, but they require supervised data and training or fine-tuning. As seen in Fig. 11 (i), large segmentation models cannot detect the transparent fin of the fish, while LPM and ours can. Moreover, LPM utilizes only one resolution, so it cannot find rocks in the river separately. As seen in Fig. 11 (ii), ODISE [49] and SAM [22] fail to segment minute object parts, like fingernails, while LPM and ours can find those segments. Furthermore, our method provides precise boundaries and segments in higher resolutions than LPM. Moreover, LPM uses Stable Diffusion [38] and requires real image inversion to find segments, while our method does not since it is based on IP2P [6]. For this reason, LPM requires more than 1 minute to perform, while our proposal takes only 10-15 seconds per image. As a result, in a direct comparison to LPM, our method has the advantage of having higher-resolution segmentation maps segmentation of more details, and it is significantly faster. The publicly available official implementations of LPM [^3] https://github.com/orpatashnik/local-prompt-mixing, SAM [^4] https://segment-anything.com/demo and ODISE [^5] https://github.com/NVlabs/ODISE are used for the results in  Fig. 11. Additionally, the same number of clusters is used for LPM and ours to achieve a fair comparison.
@@ 8.3 Qualitative comparison on segmentation maps
Figure 11: Segmentation Qualitative. Comparison between the-state-of-art segmentation methods on challenging examples.
8.4 Ablation study
In addition to the ablation study in Sec. 5.5, we also analyze token selection during cross-attention regularization as defined in Sec. 4.2. Instead of regularizing the attention of unrelated tokens, such as <start of text>, padding, and stop words, by penalizing it. We could think of doing the opposite and give high values to relevant tokens (denoted as $\tilde{S}$) within the RoI as reported in the following equation:
$\begin{split}R(QK^{T},M)=\begin{cases}QK^{T}_{ijt}\;{\color[rgb]{0.21,0.49,0.74}\definecolor[named]{pgfstrokecolor}{rgb}{0.21,0.49,0.74}\pgfsys@color@rgb@stroke{0.21}{0.49}{0.74}\pgfsys@color@rgb@fill{0.21}{0.49}{0.74}\textbf{+}}\;\alpha,&\text{if }M_{ij}=1\text{ and }t\in\tilde{S}\\ QK^{T}_{ijt},&\text{otherwise},\end{cases}\end{split}$ (4)
where $\alpha$ is a large value. This assignment guarantees that the relevant tokens related to edit instructions will have high attention scores after the softmax operation. As seen in Tab. 5, there is no significant improvement if the unrelated tokens are penalized instead of awarding the related tokens. However, penalizing the unrelated tokens gives the freedom to distribute the attention scores among relevant tokens to the process unequally. Thus, it means soft assignment among the related tokens.
@@ 8.4 Ablation study
Table 5: Ablation Study on Token Selection. For fair comparison, all parameters are the same for all settings except the ablated parameter. The top-performing is highlighted in bold, while the second-best is denoted with underline for each block.
8.5 More Qualitative Results
This section presents additional qualitative results derived from our method, emphasizing its improved effectiveness against established baselines, such as IP2P [6] and IP2P w/MB [52]. Figure 12 illustrates the application of our method in localized image editing tasks. Specifically, it demonstrates our method’s proficiency in altering the color of specific objects: (a) ottoman, (b) lamp, (c) carpet, and (d) curtain. Unlike the baseline methods, which tend to entangle the object of interest with surrounding elements, our approach achieves precise, disentangled edits. This is not achieved by the baseline that tends to alter multiple objects simultaneously rather than isolating changes to the targeted region. The disentangled and localized edits showcased in Fig. 12 highlight the potential of LIME in end-user applications where object-specific edits are crucial.
[…]
@@ 8.5 More Qualitative Results
Figure 13 demonstrates additional examples of our method’s performance on the MagicBrush [52] test set and the PIE-Bench [20] dataset. Our approach effectively executes various tasks, such as (a) replacing an animal, (b) modifying multiple objects, (c) altering the texture of an animal, and (d) changing the color of multiple objects. As illustrated in Fig. 13, our method demonstrates significant improvements over existing baselines. For instance, while baseline models like IP2P w/MB in (a) achieve reasonable edits, they often inadvertently modify areas outside the RoI, as observed in cases (b) and (c). Notably, our method helps focus the baseline models on the RoI, as seen in (b), (c), and (d), where baselines struggle to preserve the original image. Although our method is dependent on the baseline and may occasionally induce unintended changes in peripheral areas, e.g., the floor’s color, it consistently outperforms the baseline models in terms of targeted and localized editing.
Figure 14 provides further comparative analyses using the Emu-Edit test set [41]. Our method successfully handles diverse tasks, including (a) modifying parts of animals, (b) altering the color of specific objects, (c) adding, and (d) removing objects. As depicted in Fig. 14, our approach significantly surpasses existing baselines in performance. Notably, while baseline models tend to alter entire objects rather than their individual parts, e.g., animal legs, our method targets and modifies the specific sections as instructed in scenario (a). Furthermore, baseline models often inadvertently affect areas beyond the intended RoI, as seen in cases (b) and (c). In contrast, our method demonstrates precision by confining its operations within the RoI. Particularly in scenario (d), while baseline models such as IP2P struggle to maintain the integrity of the original image or, as in the case of IP2P w/MB, fail to effectively remove objects, our method excels by accurately removing the specified objects, underscoring its superiority in targeted image manipulation tasks.
9 Broader Impact & Ethical Considerations
@@
Change the alcohol shelf into a bookshelf.
Change the robin to a silver robin sculpture.
@@
Figure 13: More Qualitative Examples. We test our method on different tasks: (a) replacing an animal, (b) changing existing objects, (c) changing texture, and (d) changing the color of multiple objects. Examples are taken from established benchmarks [52, 20]. The integration of LIME enhances the performance of all models, enabling localized edits while maintaining the integrity of the remaining image areas.
Input Image
IP2P [6]
+ LIME
IP2P [6] w/MB [52]
+ LIME
Change the legs to be bionic.
Change the color of the lighthouse into completely red.
Put Stone Henge as the background of the scene.
Remove the forks from the shelf.
Figure 14: More Qualitative Examples. We test our method on different tasks: (a) editing animal parts, (b) changing color, (c) adding, and (d) removing objects. Examples are taken from established papers [41]. The integration of LIME enhances the performance of all models, enabling localized edits while maintaining the integrity of the remaining image areas.
ltx_flex_size_many"> Refer to caption
Remove the forks from the shelf.
Figure 14: More Qualitative Examples. We test our method on different tasks: (a) editing animal parts, (b) changing color, (c) adding, and (d) removing objects. Examples are taken from established papers [41]. The integration of LIME enhances the performance of all models, enabling localized edits while maintaining the integrity of the remaining image areas.

### v2

@@ A.4.1 VQGAN-CLIP
Figure 11: Investigating CLIP-T for VQGAN-CLIP [10]. CLIP-T metrics are reported below each image and calculated between the output caption and the corresponding image. Input images and edit instructions are pictured in the first column. Ground truth edit images are taken from the MagicBrush dataset.
A.4.2 Diffusion Disentanglement
Wu et al. [50] propose a disentangled attribute editing method. Since it also claims disentangled (localized) edits, we visually compare our method + IP2P with Diffusion Disentanglement. This method is not disegned for instruction-based editing, so we use input and output captions during the comparison. Figure 12 shows edit types such as (a) texture editing and (b) replacing the object with a similar one. Diffusion Disentanglement on (a) alters the background objects in the image, e.g., adding snow on and changing the shape of the branch, and also changes the features of the object of interest, e.g., removing the tail of the bird. On (b), it fails to perform the desired edit altogether. Moreover, it requires a GPU with > 48 GB RAM [^2] https://github.com/UCSB-NLP-Chang/DiffusionDisentanglement, and one image takes approximately 10 minutes on an NVIDIA A100 80GB to generate the edited version. In comparison, LIME achieves higher visual quality and takes 25 seconds to complete on NVIDIA A100 40GB with a GPU RAM usage of 25 GB.
Input Image
DiffusionDisent. [50]
Ours
Change the robin to a silver robin sculpture.
Turn the brown horse into a pink unicorn.
A.4.3 Blended Latent Diffusion
As shown in Tab. 4, Blended Latent Diffusion [2] has better results than baselines and our method. However, as shown in Fig. 13, even if their method can perform the desired edit on the RoI from the user, (a) it distorts the location of the features, e.g., heads of the birds, and (b) it loses the information of the object in the input image and creates a new object in the RoI, e.g., blanket in (b). On the other hand, our method performs visually appealing edits on the input images considering the given edit instructions by preserving as many details from the input image as possible. This is also highlighted by a significantly lower Distance metric for our method in Tab. 4.
@@ A.4.3 Blended Latent Diffusion
A.5 Qualitative comparison on segmentation maps
Our method proposes a segmentation method based on the clustering of intermediate features of the diffusion process. In this section, we provide a qualitative comparison to other segmentation methods that could be used as an alternative to this strategy. LPM [36] uses self-attention features from one resolution, such as $32\times 32$, while our method leverages the intermediate features from different resolutions to enhance the segmentation map. Then, both apply a clustering method to find the segments in the input image. Another way to find segments is by using large segmentation models, e.g., SAM [23], ODISE [51] …. As seen in Fig. 14 (i), large segmentation models cannot detect the transparent fin of the fish, while LPM and ours can. Moreover, LPM utilizes only one resolution, so it cannot find rocks in the river separately. As seen in Fig. 14 (ii), ODISE [51] and SAM [23] fail to segment minute object parts, like fingernails, while LPM and ours can find those segments. Furthermore, our method provides precise boundaries and segments in higher resolutions than LPM. Moreover, LPM uses Stable Diffusion [41] and requires real image inversion to find segments, while our method does not since it is based on IP2P [7]. For this reason, LPM requires more than 1 minute per image, while our proposal takes only 10-15 seconds per image. As a result, in a direct comparison to LPM, our method has the advantage of having higher-resolution segmentation maps with more details, and it is significantly faster. The publicly available official implementations of LPM [^3] https://github.com/orpatashnik/local-prompt-mixing, SAM [^4] https://segment-anything.com/demo and ODISE [^5] https://github.com/NVlabs/ODISE are used for the results in  Fig. 14. Additionally, the same number of clusters is used for LPM and ours to achieve a fair comparison.
@@ A.5 Qualitative comparison on segmentation maps
Figure 14: Segmentation Qualitative. Comparison between the-state-of-art segmentation methods on challenging examples. In the first example, both LPM and LIME can detect the fish fin, while SAM and ODISE cannot. Moreover, LIME provides a higher resolution than LPM. In the second example, SAM and ODISE fail to identify fingernails. Although LPM can find them, it delivers low-resolution and imprecise segments. Our method, however, accurately identifies fingernails and offers precise boundaries.
A.6 Comparison with open-vocabulary segmentors
An alternative to our proposed edit localization step, Sec. 4.1, would be to use off-the-shelves Open Vocabulary Segmentation models (OVS) and combine them with our edit application, see Sec. 4.2. A key difference between OVS and our localization method is that OVS requires an additional input, which is the object of interest to segment, and this could be a strong limitation in an instruction edit setting because the target is not always obvious from the instruction. As seen in Tab. 6, both, SEEM [56] and OV-SEG [25], significantly underperform compared to LIME.
Table 6: OVS methods for localizing the edit. We replace the edit localization part with OVS methods to show the significance of our proposed localization method. The methods are evaluated on MagicBrush dataset.
IP2P [7] 0.112 0.037 0.852 0.743 0.276
Mask Type OV-SEG [25] 0.084 0.031 0.887 0.851 0.283
SEEM [56] 0.079 0.022 0.903 0.866 0.289
IP2P + LIME 0.058 0.017 0.935 0.906 0.293
In addition to quantitative analysis in Tab. 6, we provide a qualitative comparison for localizing the edit. As seen in Fig. 15, the OVS methods do not provide precise RoI for the edit instructions even when provided with a curated additional input, e.g., object of interest. On the other hand, our proposed localization method in Sec. 4.1 provides precise and relevant RoIs without additional requirements.
Input Image
OV-SEG [25]
SEEM [56]
Ours
a
Make her outfit black.
a
Put blue glitter on fingernails.
Figure 15: OVS Method for Localization. RoI examples that are found by different OVS methods, OV-SEG [25] and SEEM [56], and our proposal in Sec. 4.1. Since OVS methods require object of interest, words with bold are used for the object of interest.
A.7 Ablation study
A.7.1 Related Token Rewarding
In addition to the ablation study in Sec. 5.5, we also analyze token selection during cross-attention regularization as defined in Sec. 4.2. Instead of regularizing the attention of unrelated tokens, such as <start of text>, padding, and stop words, by penalizing it, we could think of doing the opposite and give high values to relevant tokens (denoted as $\tilde{S}$) within the RoI as reported in the following equation:
$\begin{split}R(QK^{T},M)=\begin{cases}QK^{T}_{ijt}\;{\color[rgb]{0.12,0.49,0.85}\definecolor[named]{pgfstrokecolor}{rgb}{0.12,0.49,0.85}\textbf{+}}\;\infty,&\text{if }M_{ij}=1\text{ and }t\in\tilde{S}\\ QK^{T}_{ijt},&\text{otherwise},\end{cases}\end{split}$ (4)
This assignment guarantees that relevant tokens related to edit instructions have high scores after the softmax operation. As seen in Tab. 7, there is no significant improvement if the unrelated tokens are penalized instead of awarding the related tokens. However, penalizing the unrelated tokens gives the freedom to distribute the attention scores unequally among relevant tokens. Thus, it allows for a soft assignment of areas of the image among the related tokens.
Table 7: Ablation Study on Token Selection. For fair comparison, all parameters are the same for all settings except the ablated parameter.
@@ A.7.1 Related Token Rewarding
In addition to quantitative analysis on the MagicBrush dataset, Fig. 16 shows the attention scores for rewarding related tokens vs regularizing unrelated tokens. Even though the final edit results and the last column are not significantly different, unrelated token regularization results in uneven attention scores among related tokens, while related token regularization leads to equal attention scores.
Token-based cross-attention probabilities
Unrelated
Related
Figure 16: Ablation study on related token rewarding. Instead of regularizing the unrelated tokens, see the first row, we ablate the opposite which is rewarding the related tokens by giving + $\infty$ within RoI, see the second row.
A.7.2 Clustering Method Alternatives
Since our implementation uses KMeans, it requires a number of clusters parameter to be decided and fixed. While in the main paper, we show that a good tuning of this hyperparameter works robustly across all the evaluated datasets, in this section, we perform a preliminary exploration of alternative clustering techniques that would relax this assumption. We replace K-means with Agglomerative Clustering, which does not require this parameter. For this implementation, we use the cosine similarity metric between features since [46] shows that the cosine similarity provides significant information about semantics. We provide a minimum distance threshold, which can be between 0-1, and it automatically determines the number of clusters. The higher the threshold, the more clusters. As seen in Tab. 8, this variation achieves similar performance by properly tuning the distance threshold. This ablation study proves that our method is robust for selecting the clustering method, and any clustering method can be combined with LIME.
Table 8: Ablation Study on Clustering Method. For all experiments, IP2P is the base architecture and the evaluation is on the MagicBrush dataset. Each parameter is modified separately, while other parameters are kept fixed to isolate their impact.
IP2P [7] 0.112 0.037 0.852 0.743 0.276
# of Clusters $4$ 0.080 0.022 0.923 0.885 0.295
$8$ 0.058 0.017 0.935 0.906 0.293
$16$ 0.062 0.018 0.933 0.903 0.294
$32$ 0.064 0.018 0.932 0.901 0.291
Distance Threshold $0.7$ 0.080 0.022 0.927 0.893 0.294
$0.6$ 0.076 0.021 0.929 0.894 0.294
$0.5$ 0.063 0.018 0.933 0.902 0.293
$0.4$ 0.072 0.020 0.930 0.896 0.293
Input Image
@@ A.7.2 Clustering Method Alternatives
A.8 More Qualitative Results
This section presents additional qualitative results derived from our method, emphasizing its improved effectiveness against established baselines, such as IP2P [7] and IP2P w/MB [53]. Figure 17 illustrates the application of our method in localized image editing tasks. Specifically, it demonstrates our method’s proficiency in altering the color of specific objects: (a) ottoman, (b) lamp, (c) carpet, and (d) curtain. Unlike the baseline methods, which tend to entangle the object of interest with surrounding elements, our approach achieves precise, disentangled edits. This is not achieved by the baseline, which tends to alter multiple objects simultaneously rather than isolating changes to the targeted region. The disentangled and localized edits showcased in Fig. 17 highlight the potential of LIME in end-user applications where object-specific edits are crucial.
Figure 19 demonstrates additional examples of our method’s performance on the Emu-Edit test set [44], the MagicBrush [53] test set and the PIE-Bench [21] dataset. Our approach effectively executes various tasks, such as (a) replacing an animal, (b) modifying parts of animals, and (c) changing the color of multiple objects. As illustrated in Fig. 19, our method demonstrates significant improvements over existing baselines. For instance, while baseline models like IP2P w/MB in (a) achieve reasonable edits, they often inadvertently modify areas outside the RoI, as observed in cases (b) and (c). Notably, our method helps focus the baseline models on the RoI, as seen in (b) and (c), where baselines struggle to preserve the original image. Although our method is dependent on the baseline and may occasionally induce unintended changes in peripheral areas, e.g., the floor’s color, it consistently outperforms the baseline models in terms of targeted and localized editing.
B Implementation Details
We obtain the results using an NVIDIA A100 40GB GPU machine. For $512\times 512$ images the IP2P-based baselines (e.g., IP2P, IP2P w/MB, HIVE, and HIVE w/MB) take approximately 15 seconds per edit, while for LIME integrated models, it takes $\approx$25 seconds.
B.1 User Study Setting
We carry out a study with 54 questions involving users, asking 53 anonymous individuals on the crowd-sourcing platform Prolific [38]. In our user study, participants will be presented with two alternative edited images alongside their corresponding input images and editing instructions. They will be tasked with evaluating the effectiveness of the edits in achieving the specified outcome and the ability of the editing method to preserve the details in areas not targeted by the instruction. Using the example provided in Fig. 18, where the editing instruction is to Change to a rosé, participants must discern which edited image (a, b or neither) not only best satisfies this directive but also maintains the fidelity of the scene’s irrelevant aspects. The aggregated data from participant responses will yield insights into the preferred methods for both accurate editing and detail preservation, thereby influencing the development of advanced image editing methods.
Figure 18: User study setting. The example with the edit instruction with the corresponding input image and randomly ordered two edited images from baseline and baseline + LIME.
@@ B.1 User Study Setting
Change the legs to be bionic.
@@
Figure 19: More Qualitative Examples. We test our method on different tasks: (a) replacing an animal, (b) editing animal parts, and (c) changing the color of multiple objects. The integration of LIME enhances the performance of all models, enabling localized edits while maintaining the integrity of the remaining image areas.
B.2 Reproducibility Statement
We guarantee that all the results presented in the main manuscript and the supplementary materials can be reproduced. We will make our code base containing examples and hyperparameters public so that the results can be reproduced.
B.3 Baselines
Open-Edit [27]: This GAN-based approach uses a reconstruction loss for pre-training and incorporates a consistency loss during fine-tuning on specific images. Its unique feature is the arithmetic manipulation of word embeddings within a shared space of visual and textual features.
VQGAN-CLIP [10]: Enhancing the VQGAN [13] framework with CLIP embeddings [39], this method fine-tunes VQGAN using the similarity of CLIP embeddings between the generated image and the target text, leading to optimized image generation.
SDEdit [29]: Leveraging the capabilities of Stable Diffusion [41], SDEdit introduces a tuning-free approach. It uses stochastic differential equation noise, adding it to the source image and subsequently denoising to approximate the target image, all based on the target caption.
Text2LIVE [4]: It propose a Vision Transformer [12] for generating edited objects on an additional layer. It incorporates data augmentation and CLIP [39] supervision, ultimately alpha-blending the edited layer with the original to create the target image.
Null Text Inversion [33]: By optimizing the DDIM [45] trajectory, this method initially inverts the source image. After, it performs image editing during the denoising process guided by cross-attention [18] between text and image.
SINE [55]: Real images are edited using model-based guidance and patch-based fine-tuning process.
DreamBooth [42]: It fine-tunes a diffusion model by learning special text tokens and adjusting model parameters on a set of images for editing.
Textual-Inversion [15]: It fine-tunes a token embedding within the text-encoder space using a set of images.
Imagic [22]: It edits images through a three-step process: first fine-tuning a token embedding, then fine-tuning the parameters of a text-guided image diffusion model using the fine-tuned token embedding, and finally performing interpolation to generate various edits based on a target prompt.
DiffEdit [9]: It identifies the region to edit in images by contrasting between a conditional and unconditional diffusion model based on query and reference texts. Then, it reconstructs the edited image by collecting the features from the text-query by combining the features in the noise/latent space, considering the region to edit.
Blended Latent Diffusion [2]: This method uses a text-to-image Latent Diffusion Model (LDM) to edit the user-defined mask region. It extracts features for the mask region from the edit text, and for the rest of the image, it uses features from the original image in the noise/latent space.
DirectDiffusion [21]: It inverts the input image into the latent space of Stable Diffusion [41] and then applies Prompt2Prompt [18] to obtain the desired edit without making any changes to the edit diffusion branch.
Diffusion Disentanglement [50]: It finds the linear combination of the text embeddings of the input caption and the desired edit to be performed. Since it does not fine-tune Stable Diffusion parameters, they claim that the method performs disentangled edits.
InstructPix2Pix (IP2P) [7]: Starting from the foundation of Stable Diffusion [41], the model is fine-tuned for instruction-based editing tasks. It ensures that the edited image closely follows the given instructions while maintaining the source image without the need for test-time tuning.
InstructPix2Pix w/MagicBrush [53]: A version of IP2P [7] trained on MagicBrush train set [53]. Since the MagicBrush dataset has more localized edit examples, the fine-tuned version has better results, as seen in Fig. 6(a).
HIVE [54]: It extends IP2P [7] by fine-tuning it with an expanded dataset. Further refinement is achieved through fine-tuning with a reward model, which is developed based on human-ranked data.
HIVE w/MagicBrush [53]: HIVE [54] fine-tuned on MagicBrush train set [53]. Since the MagicBrush dataset has more localized edit examples, the fine-tuned version has better results, as seen in Fig. 6(a).
C Broader Impact & Ethical Considerations